OpenStudy (starryknight):

please explain in steps please :)))) Use the technique of completing the square to transform the quadratic equation below into the form (x + c)2 = a. 4x2 + 16x + 12 = 0 answer choices: (x + 4)2 = -1 (x + 16)2 = -1 (x + 16)2 = 1 (x + 2)2 = 1

OpenStudy (anonymous):

first factor out a 4 to get \[4(x^2+4+3)=0\] divide both sides by 4 to get \[x^2+4+3=0\] then add 1 to both sides to get \[x^2+4x+4=1\] then you have a perfect square.. \[(x+2)^2=1\]

OpenStudy (starryknight):

thank you :)))

OpenStudy (anonymous):

you are welcome:)